OpenStudy (anonymous):

Evaluate the indefinite integral. 1/cos^5(x) dx

OpenStudy (kira_yamato):

\[∫\frac{1}{\cos^5x}dx=∫(sec^5 x dx)\] Let w = sec x, so sec^5 x = w^5/2 dw/dx = sec x tanx dw = sec x tan x dx \[dw = 2w \sqrt{w^2 - 1}dx\] Do integration by substitution from here

OpenStudy (anonymous):

oh okay thanks for the help

OpenStudy (anonymous):

is dw the same with du?!

OpenStudy (kira_yamato):

Yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (jamesj):

No, this isn't a helpful substitution, since you will nave to reintroduce trig functions to get any further with it. Instead try this: Apply integration by parts, stripping out a sec^2 term and integrating it to tan x, to find show: \[\int\limits \sec^n x \ dx = {1 \over n-1} \sec^{n-2} x \tan x + {n-2 \over n-1} \int\limits \sec^{n-2} x \ dx\] Then apply it to your integrand, sec^5 x, twice.